OpenStudy (anonymous):

Let f(x) = x^2 – 4x – 7 and g(x) = x – 3. Find f(g(x)). Can someone please explain this to me?, I really don't understand how to do these kinds of functions.

OpenStudy (anonymous):

Alright,this is actually what is mathematically referred to as Function composition ( http://en.wikipedia.org/wiki/Function_composition) Now,in a very simple term,which is sufficient for this problem,to find f(g(x)), take the function of g(x) and put it into f(x) for all x's, so here, \[ f(g(x))= f((x – 3))= (x – 3)^2 - 4 (x – 3) -7 = 14-10 x+x^2 \]